Write the following rectangular points in polar coordinates. i. (0.1) ii. (5,5) 333 iii. 21 2 iv. (-6, 0) v. (9, -3)
Answer #1

0 We have give rectangular Gor points and we have to change rectangular points to polar Coordinates. To Convert from reed angso 150 ET? ty? - tis)+(5)2 = or tuman Box) = tanil () tastro) Vly so (5,5) :) → (150, 45) (5,5) ( 7.071, (- 315 - x2 + y = TC) (-6,0) na r6Iy=0 = Jer2+6012 6 No ² ty? fan-(7) tam () = tant (0) = 0 (6,0) so (-6,0) (-6,0) (6,0) du V (9,-3) n=9, y=-3